# Achieving Collective Welfare in Multi-Agent Reinforcement Learning via Suggestion Sharing

### v1

@@ Abstract
In human society, the conflict between self-interest and collective well-being often obstructs efforts to achieve shared welfare. Related concepts like the Tragedy of the Commons and Social Dilemmas frequently manifest in our daily lives. As artificial agents increasingly serve as autonomous proxies for humans, we propose using multi-agent reinforcement learning (MARL) to address this issue—learning policies to maximise collective returns even when individual agents’ interests conflict with the collective one. Traditional MARL solutions involve sharing rewards, values, and policies or designing intrinsic rewards to encourage agents to learn collectively optimal policies. We introduce a novel MARL approach based on Suggestion Sharing (SS), where agents exchange only action suggestions. This method enables effective cooperation without the need to design intrinsic rewards, achieving strong performance while revealing less private information compared to sharing rewards, values, or policies. Our theoretical analysis establishes a bound on the discrepancy between collective and individual objectives, demonstrating how sharing suggestions can align agents’ behaviours with the collective objective. Experimental results demonstrate that SS performs competitively with baselines that rely on value or policy sharing or intrinsic rewards.
@@ 1 Introduction
Multi-agent reinforcement learning enables collaborative decision-making in diverse real-world applications, such as autonomous vehicle control [Xia et al., 2022, Qiu et al., 2023, Jin et al., 2021], robotics [Wang et al., 2022, Peng et al., 2021, Sun et al., 2020], and communications systems [Siedler and Alpha, , Huang and Zhou, 2022]. In these scenarios, artificial agents often act as autonomous decision makers. MARL provides a powerful framework for these settings, enabling agents to learn coordination strategies based on rewards reflecting a common goal.
@@
In practice, agents typically do not have access to others’ rewards, value functions, or policy functions. For instance, in a smart grid system, consumers’ electricity usage policies reflect sensitive information about their daily routines and financial constraints, and their interests (rewards/ values) related to comfort are also private. This information may not be something they are willing to share with other participants or a central coordinator. Similarly, an a traffic network, drivers’ routing preferences and time valuations, which reveal sensitive details about their destinations, schedule constraints, and willingness to pay for faster travel, are rarely shared with others. Traditional MARL approaches that rely on agents sharing rewards, policies, or value estimates thus become problematic in such settings.
@@
The main contributions of this paper are as follows. We propose using MARL to address sociological problems such as Social Dilemmas and the Tragedy of the Commons, in scenarios where agents can make decisions or provide advice to humans while communicating within systems like smart grids and cooperative driving systems. We introduce the SS algorithm, which promotes cooperation for collective welfare while reducing the sharing of private information. Theoretically, we show that the optimisation objective of SS serves as a lower bound for the original collective objective. Empirical results demonstrate that SS performs competitively with existing MARL algorithms that rely on sharing policies or values.
@@ 3 Preliminaries and Problem Statement
$\footnotesize\eta(\boldsymbol{\pi})=\sum_{i=1}^{N}\mathbb{E}_{\tau\sim\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}^{i}\right],$ (1)
@@
$\footnotesize A_{i}^{\boldsymbol{\pi}}(s,\boldsymbol{a})=Q_{i}^{\boldsymbol{\pi}}(s,\boldsymbol{a})-V_{i}^{\boldsymbol{\pi}}(s)$ (2)
@@
However, in our setup, agents do not have direct access to others’ policies, rewards, or values. This setting is particularly relevant for applications where users prefer not to reveal their exact policies and rewards or values. Our work aims to bridge this gap between individual and collective return maximisation. It enables agents to approximate the optimisation of the collective objective while operating solely with their individual reward signals. In the next section, we present a method where agents iteratively share suggestions to maximise a lower bound of Eq.1. This method is general and not dependent on any specific protocol for communicating suggestions between agents. In Sec.4.3, we propose a practical algorithm that involves sharing information within agents’ neighbourhoods. Our experiments demonstrate the effects of different sharing protocols on the performance of MARL cooperation.
@@ 4.1 Theoretical developments
$\footnotesize\zeta_{\boldsymbol{\pi}^{\prime}}(\boldsymbol{\tilde{\Pi}})=\sum_{i}\mathbb{E}_{s\sim d^{\boldsymbol{\pi}^{\prime}}(s),\boldsymbol{a}\sim\boldsymbol{\tilde{\pi}}^{i}(\boldsymbol{a}|s)}\left[A_{i}^{\boldsymbol{\pi}^{\prime}}(s,\boldsymbol{a})\right],$ (6)
@@
$\footnotesize f^{\boldsymbol{\pi}^{\prime}}=\sum_{i}\frac{1}{2}\max_{s,\boldsymbol{a}}\left|A_{i}^{\boldsymbol{\pi}^{\prime}}(s,\boldsymbol{a})\right|\cdot|\mathcal{A}|\cdot\|d^{\boldsymbol{\pi}^{\prime}}\|_{2}^{2},$ (8)
and $\|d^{\pi^{\prime}}\|_{2}^{2}=\sum_{s}(d^{\pi^{\prime}}(s))^{2}$.
@@
$\footnotesize\zeta_{\boldsymbol{\pi}_{old}}(\boldsymbol{\tilde{\Pi}}_{new})=\sum_{i}\mathbb{E}_{s\sim d^{\boldsymbol{\pi}_{old}}(s),\boldsymbol{a}\sim\boldsymbol{\tilde{\pi}}_{new}^{i}(\boldsymbol{a}|s)}\left[A_{i}^{\boldsymbol{\pi}_{old}}(s,\boldsymbol{a})\right],$ (9)
@@ 4.2 A Surrogate Optimisation Objective
To leverage this insight, we use the lower bound given by Theorem 1. By maximising this lower bound , we can maximise the collective return. We can ignore the terms $\eta(\boldsymbol{\pi}_{\text{old}})$ and $f^{\boldsymbol{\pi}_{\text{old}}}$ from Theorem 1 in our optimisation problem, as they are not relevant to optimising $\boldsymbol{\tilde{\Pi}}$ and their values are usually bounded. To be specific, the value of $\eta(\boldsymbol{\pi}_{\text{old}})$ is bounded as the reward value is bounded. For $f^{\boldsymbol{\pi}_{\text{old}}}$, as defined in Eq. 8, its value is also bounded since (1) We focus on scenarios with finite and relatively small action spaces, which are common in many real-world applications, so $|\mathcal{A}|$ (the size of the action space) is not excessively large. (2) The term $\|d^{\pi_{\text{old}}}\|^{2}_{2}$ is the square L2-norm of the state visitation distribution, which is bounded.(3) The advantage function $A^{\pi_{\text{old}}}_{i}(s,a)$ is also bounded as the reward value is bounded.
@@
The constraints in Eq. 12 are imposed on $\pi^{ii}$ and $\pi^{ij}$ ($j\neq i$), which together compose $\boldsymbol{\tilde{\pi}}^{i}$. Therefore, these constraints effectively limit the space of possible $\boldsymbol{\tilde{\pi}}^{i}$ by constraining its components. Constraint (a) limits how much the agent’s own policy can change, while constraints (b) and (c) ensure that the suggestions are close to the actual policies of other agents. The constraints also depend on other agents’ policies $\pi^{jj}$ and their suggestions for agent $i$’s policy, $\pi^{ji}$. To enable the evaluation of these terms, each agent $j$ shares its action distribution $\pi^{jj}(\cdot|s)$ and the action distribution suggestion $\pi^{ji}(\cdot|s)$ with agent $i$. This sharing enables each agent $i$ to assess the constraint terms, which couples individual advantage optimisations under local constraints. These constraints reflect both the differences between the policies of others and an agent’s suggestions on them, as well as the discrepancy between an agent’s own policy and others’ suggestions on it. By distributing the optimisation while exchanging policy suggestions, this approach balances individual policy updates while maintaining global coordination among agents.
@@
4.3 A Practical Algorithm for Learning with SS
@@
$\footnotesize\xi_{i}=\frac{\pi^{ii}(a_{i}|s^{\prime};\theta^{ii})}{\pi^{ii}_{old}(a_{i}|s^{\prime};\theta^{ii}_{old})},\quad\xi_{\mathcal{N}_{i}}=\prod_{j\in\mathcal{N}_{i}}\frac{\pi^{ij}(a_{j}|s;\theta^{ij})}{\pi^{jj}_{old}(a_{j}|s;\theta^{jj}_{old})}.$ (13)
@@
$\footnotesize\mathbb{I}_{X}(s,\boldsymbol{a})=\begin{cases}1&\text{if }(s,\boldsymbol{a})\in X,\\ 0&\text{otherwise}.\end{cases}$ (15)
@@
where $\theta^{ii}$ denotes the parameters of $\pi^{ii}$ and $\boldsymbol{\theta}^{-ii}$ denotes the parameters of all the $\pi^{ij}$ ($j\in\mathcal{N}i$). With this objective, each agent optimises its own policy $\pi^{ii}$ under the constraint of staying close to the suggested policies. In the meanwhile, the suggestions $\pi^{ij}$ which are involved in $\xi_{\mathcal{N}_{i}}$, are optimised to maximise the agent’s individual advantage function $A_{i}$ under the constraint of avoiding deviating too far from the actual policies of other agents. This objective function balances individual policy updates with the need for coordination among agents, thereby aligning individual objectives with collective goals. Fig. 6 in Appendix A.2 shows an illustration of our method.
4.4 Implementation Details
@@
$\footnotesize\mathcal{L}_{V_{i}}=\mathbb{E}[(V_{i}(s_{t})-\sum_{l=0}^{\infty}\gamma^{l}r_{i}^{t+l})^{2}].$ (18)
Algorithm 1 in Appendix A.2 presents the detailed procedure used in our experimental section.
@@ 5.1 Environments
Cooperative predation (C. Predation). This environment involves a sequential social dilemma in a continuous domain, where multiple predator agents aim to capture a single prey. All predators cooperating (approaching the prey) results in each receiving a reward of $-1$. Universal defection (not approaching) yields a reward of $-2N+1$ for each predator, where $N$ is the total number of agents. In mixed scenarios, predators pursuing the prey receive a reward of $-2N$, while non-participating predators gain 00. The challenge is to incentivise agents to cooperate and capture the prey rather than acting selfishly. At the start of each episode, the prey’s position, $x_{tar}\in\mathcal{X}$, and the agents’ initial positions, $x_{ag_{i}}\in\mathcal{X}$, are randomly assigned within $\mathcal{X}=[0,30]$. The state is represented as $s^{t}=[x_{ag_{1}}^{t}-x_{tar},\ldots,x_{ag_{N}}^{t}-x_{tar}]$, a continuous variable. The action set $\mathcal{A}={-1,+1}$ corresponds to left and right movements. Neighbouring agents are defined as those within a normalised distance of $0.1$. Fig. 7(b) in Appendix A.3 illustrates this environment. The episode time horizon is set to $30$, and our main experiments use $8$ predator agents.
@@ 5.3 Experimental Results
Main results. We conducted 5 runs with different seeds for each algorithm and environment. Fig. 1 shows the training curves and Fig.  2 the normalised final averaged returns for different algorithms. The averaged return refers to the collective return, normalised by the number of agents and episode length. Our SS algorithm demonstrates consistently strong performance across all tasks, with averaged returns matching or exceeding those of baseline algorithms that rely on sharing values or policy parameters. This highlights the effectiveness of SS.
In Fig. 1, SS converges faster than PS, which implies that sharing action distributions is more efficient than sharing parameters of policy networks. In Fig. 2, SS outperforms both VS and VPS. Additionally, PS shows better performance than VS and VPS, which may indicate that sharing policy information is more effective than sharing value information. Notably, SS outperforms CL in some cases. We hypothesise that in these scenarios, SS facilitates cooperation by enabling agents to encourage each other through action suggestions based on their individual interests, while CL may struggle due to exploration issues arising from a lack of successful cooperation experiences. IMR also shows competitive performance, even achieving the best results in a specific case. However, for the problem addressed in this paper, adding intrinsic rewards may not always be practical, especially in scenarios where designing appropriate intrinsic rewards is challenging.

### v2

@@ Abstract
In human society, the conflict between self-interest and collective well-being often obstructs efforts to achieve shared welfare. Related concepts like the Tragedy of the Commons and Social Dilemmas frequently manifest in our daily lives. As artificial agents increasingly serve as autonomous proxies for humans, we propose a novel multi-agent reinforcement learning (MARL) method to address this issue—learning policies to maximise collective returns even when individual agents’ interests conflict with the collective one. Unlike traditional cooperative MARL solutions that involve sharing rewards, values, and policies or designing intrinsic rewards to encourage agents to learn collectively optimal policies, we propose a novel MARL approach where agents exchange action suggestions. Our method reveals less private information compared to sharing rewards, values, or policies, while enabling effective cooperation without the need to design intrinsic rewards. Our algorithm is supported by our theoretical analysis that establishes a bound on the discrepancy between collective and individual objectives, demonstrating how sharing suggestions can align agents’ behaviours with the collective objective. Experimental results demonstrate that our algorithm performs competitively with baselines that rely on value or policy sharing or intrinsic rewards.
@@ 1 Introduction
Multi-agent reinforcement learning (MARL) enables collaborative decision-making in diverse real-world applications, such as autonomous vehicle control [Xia et al., 2022, Qiu et al., 2023, Jin et al., 2021], robotics [Wang et al., 2022, Peng et al., 2021, Sun et al., 2020], and communications systems [Siedler and Alpha, , Huang and Zhou, 2022]. In these scenarios, artificial agents often act as autonomous decision makers. MARL provides a powerful framework for these settings, enabling agents to learn coordination strategies based on rewards reflecting a common goal.
@@
In practice, agents typically do not have access to others’ rewards, value functions, or policy functions. For instance, in a smart grid system, consumers’ electricity usage policies reflect sensitive information about their daily routines and financial constraints, and their interests (rewards/ values) related to comfort are also private. This information may not be something they are willing to share with other participants or a central coordinator. Similarly, in a traffic network, drivers’ routing preferences and time valuations, which reveal sensitive details about their destinations, schedule constraints, and willingness to pay for faster travel, are rarely shared with others. Traditional MARL approaches that rely on agents sharing rewards, policies, or value estimates thus become problematic in such settings.
@@
The main contributions of this paper are as follows. We propose a novel Suggestion-Sharing-based MARL (SS) method to learn cooperative policies for collective welfare when individual interests may conflict with collective objectives. Our method reveals less private information than the traditional cooperative MARL methods that resort to sharing rewards, values, or policies, while enabling effective cooperation without the need to design intrinsic rewards. Theoretically, we show that the optimisation objective of SS serves as a lower bound for the original collective objective. Empirical results demonstrate that SS performs competitively with existing MARL algorithms that rely on sharing policies or values.
@@ 3 Preliminaries and Problem Statement
$\eta(\boldsymbol{\pi})=\sum_{i=1}^{N}\mathbb{E}_{\tau\sim\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}^{i}\right],$ (1)
@@
$A_{i}^{\boldsymbol{\pi}}(s,\boldsymbol{a})=Q_{i}^{\boldsymbol{\pi}}(s,\boldsymbol{a})-V_{i}^{\boldsymbol{\pi}}(s)$ (2)
@@
However, in our setup, agents do not have direct access to others’ policies, rewards, or values. This setting is particularly relevant for applications where users prefer not to reveal their exact policies and rewards or values. Our work aims to bridge this gap between individual and collective return maximisation. It enables agents to approximate the optimisation of the collective objective while operating solely with their individual reward signals. In the next section, we present a method where agents iteratively share suggestions to maximise a lower bound of Eq.1. This method is general and not dependent on any specific protocol for communicating suggestions between agents. In Sec.4.3, we propose a practical algorithm that involves sharing information within agents’ neighbourhoods. Our experiments demonstrate the effects of different sharing protocols on the performance of MARL cooperation. For convenience, notations frequently used in this paper are listed in Table 1.
Table 1: Notations frequently used in this paper.
$\eta$ The expectation of collective cumulative return of all agents
$\pi^{i}$ Individual policy
$\boldsymbol{\pi}$ Joint policy
$\pi^{ij}$ Suggestion of agent $i$ about agent $j$’s policy when $j\neq i$
$\pi^{ii}$ Equivalent to $\pi^{i}$, which is agent $i$’s own policy
$\boldsymbol{\tilde{\pi}}^{i}$ Suggesting joint policy of agent $i$
$\boldsymbol{\tilde{\Pi}}$ Collection of all suggesting joint policies across agents, i.e., $(\boldsymbol{\tilde{\pi}}^{1},\cdots,\boldsymbol{\tilde{\pi}}^{i},\cdots,\boldsymbol{\tilde{\pi}}^{N})$
$A_{i}^{\boldsymbol{\pi}}$ Individual advantage function under joint policy $\boldsymbol{\pi}$
$\hat{A}_{i}$ Estimated value of individual advantage function
$\theta^{ij}$ Parameters of $\pi^{ij}$
$\mathcal{N}_{i}$ Agent $i$’s neighbours
$\boldsymbol{\theta}^{-ii}$ Parameters of all the $\pi^{ij}$ ($j\in\mathcal{N}_{i}$)
@@ 4.1 Theoretical Developments
$\zeta_{\boldsymbol{\pi}^{\prime}}(\boldsymbol{\tilde{\Pi}})=\sum_{i}\mathbb{E}_{s\sim d^{\boldsymbol{\pi}^{\prime}}(s),\boldsymbol{a}\sim\boldsymbol{\tilde{\pi}}^{i}(\boldsymbol{a}|s)}\left[A_{i}^{\boldsymbol{\pi}^{\prime}}(s,\boldsymbol{a})\right],$ (6)
@@
and $\|d^{\pi^{\prime}}\|_{2}^{2}:=\sum_{s}(d^{\pi^{\prime}}(s))^{2}$.
@@
$\zeta_{\boldsymbol{\pi}_{old}}(\boldsymbol{\tilde{\Pi}}_{new})=\sum_{i}\mathbb{E}_{s\sim d^{\boldsymbol{\pi}_{old}}(s),\boldsymbol{a}\sim\boldsymbol{\tilde{\pi}}_{new}^{i}(\boldsymbol{a}|s)}\left[A_{i}^{\boldsymbol{\pi}_{old}}(s,\boldsymbol{a})\right],$ (9)
@@ 4.2 A Surrogate Optimisation Objective
To leverage this insight, we use the lower bound given by Theorem 1. By maximising this lower bound , we can maximise the collective return. We can ignore the terms $\eta(\boldsymbol{\pi}_{\text{old}})$ and $f^{\boldsymbol{\pi}_{\text{old}}}$ from Theorem 1 in our optimisation problem, as they are not relevant to optimising $\boldsymbol{\tilde{\Pi}}$ and their values are usually bounded. To be specific, the value of $\eta(\boldsymbol{\pi}_{\text{old}})$ is bounded as the reward value is bounded. For $f^{\boldsymbol{\pi}_{\text{old}}}$, as defined in Eq. 8, its value is also bounded since (1) We focus on scenarios with finite and relatively small action spaces (each agent’s discrete action set typically consists of 2–10 actions), which are common in many real-world applications, so $|\mathcal{A}|$ (the size of the action space) is not excessively large. (2) The term $\|d^{\pi_{\text{old}}}\|^{2}_{2}$ is the square L2-norm of the state visitation distribution, which is bounded.(3) The advantage function $A^{\boldsymbol{\pi}_{\text{old}}}_{i}(s,a)$ is also bounded as the reward value is bounded.
@@
The constraints in Eq. 12 are imposed on $\pi^{ii}$ and $\pi^{ij}$ ($j\neq i$), which together compose $\boldsymbol{\tilde{\pi}}^{i}$. Therefore, these constraints effectively limit the space of possible $\boldsymbol{\tilde{\pi}}^{i}$ by constraining its components. Constraint (a) limits how much the agent’s own policy can change, while constraints (b) and (c) ensure that the suggestions are close to the actual policies of other agents. The corresponding terms in these constrains are bounded by some constants or functions, so that they can remain finite. This boundedness aims to guarantee that the discrepancy between the collective and individual objectives is controllable.
The constraints also depend on other agents’ policies $\pi^{jj}$ and their suggestions for agent $i$’s policy, $\pi^{ji}$. To enable the evaluation of these terms, each agent $j$ shares its action distribution $\pi^{jj}(\cdot|s)$ and the action distribution suggestion $\pi^{ji}(\cdot|s)$ with agent $i$. This sharing enables each agent $i$ to assess the constraint terms, which couples individual advantage optimisations under local constraints. These constraints reflect both the differences between the policies of others and an agent’s suggestions on them, as well as the discrepancy between an agent’s own policy and others’ suggestions on it. By distributing the optimisation while exchanging policy suggestions, this approach balances individual policy updates while maintaining global coordination among agents.
@@
4.3 A Practical Algorithm for MARL with SS
@@
$\xi_{i}=\frac{\pi^{ii}(a_{i}|s^{\prime};\theta^{ii})}{\pi^{ii}_{old}(a_{i}|s^{\prime};\theta^{ii}_{old})},\quad\xi_{\mathcal{N}_{i}}=\prod_{j\in\mathcal{N}_{i}}\frac{\pi^{ij}(a_{j}|s;\theta^{ij})}{\pi^{jj}_{old}(a_{j}|s;\theta^{jj}_{old})}.$ (13)
@@
$\mathbb{I}_{X}(s,\boldsymbol{a})=\begin{cases}1&\text{if }(s,\boldsymbol{a})\in X,\\ 0&\text{otherwise}.\end{cases}$ (15)
@@
where $\theta^{ii}$ denotes the parameters of $\pi^{ii}$ and $\boldsymbol{\theta}^{-ii}$ denotes the parameters of all the $\pi^{ij}$ ($j\in\mathcal{N}_{i}$). With this objective, each agent optimises its own policy $\pi^{ii}$ under the constraint of staying close to the suggested policies. In the meanwhile, the suggestions $\pi^{ij}$ which are involved in $\xi_{\mathcal{N}_{i}}$, are optimised to maximise the agent’s individual advantage function $A_{i}$ under the constraint of avoiding deviating too far from the actual policies of other agents. This objective function balances individual policy updates with the need for coordination among agents, thereby aligning individual objectives with collective goals.
@@
$\mathcal{L}_{V_{i}}=\mathbb{E}[(V_{i}(s_{t})-\sum_{l=0}^{\infty}\gamma^{l}r_{i}^{t+l})^{2}].$ (18)
Algorithm 1 presents the detailed procedure of our practical algorithm. A corresponding illustration figure can be found in Fig. 6 in Appendix A.2.
Algorithm 1 Suggestion-Sharing-based MARL (SS)
  for episode = 1 to $E$ do
     $\mathcal{D}_{i}\leftarrow\phi,\forall i$
     Observe initial state $s_{1}$
     # Interact with the environment
     for $t=1$ to $T$ do
        Execute action $a_{t}^{i}\in\pi^{ii}(\cdot|s_{t})$
        Observe reward $r_{t}^{i}$ and next state $s_{t+1}$
        Store $(s_{t},a_{t}^{i},r_{t}^{i},s_{t+1})\in\mathcal{D}_{i}$
     end for
     for iteration = 1 to $K$ do
        # Share with agents
        for each agent $i$ do
        end for
        for $i=1$ to $N$ do
           # Learn policy and value individually
           Update $\boldsymbol{\tilde{\pi}}^{i}$ using Eq 16
           Update $V_{i}$ using Eq 18
           $\boldsymbol{\tilde{\pi}}^{i}_{old}\leftarrow\boldsymbol{\tilde{\pi}}^{i}$
           # Update sharing with agents
        end for
     end for
  end for
@@ 5.1 Environments
Cooperative predation (C. Predation). This environment involves a sequential social dilemma in a continuous domain, where multiple predator agents aim to capture a single prey. All predators cooperating (approaching the prey) results in each receiving a reward of $-1$. Universal defection (not approaching) yields a reward of $-2N+1$ for each predator, where $N$ is the total number of agents. In mixed scenarios, predators pursuing the prey receive a reward of $-2N$, while non-participating predators gain 00. The challenge is to incentivise agents to cooperate and capture the prey rather than acting selfishly. At the start of each episode, the prey’s position, $x_{tar}\in\mathcal{X}$, and the agents’ initial positions, $x_{ag_{i}}\in\mathcal{X}$, are randomly assigned within $\mathcal{X}=[0,30]$. The state is represented as $s^{t}=[x_{ag_{1}}^{t}-x_{tar},\ldots,x_{ag_{N}}^{t}-x_{tar}]$, a continuous variable. The action set $\mathcal{A}=\{-1,+1\}$ corresponds to left and right movements. Neighbouring agents are defined as those within a normalised distance of $0.1$. Fig. 7(b) in Appendix A.3 illustrates this environment. The episode time horizon is set to $30$, and our main experiments use $8$ predator agents.
@@ 5.3 Experimental Results
Main results. We conducted 5 runs with different seeds for each algorithm and environment. Fig. 1 shows the training curves and Fig.  2 the normalised final averaged returns for different algorithms. The averaged return refers to the collective return, normalised by the number of agents and episode length. Our SS algorithm demonstrates consistently strong performance across all tasks, with averaged returns matching or exceeding those of baseline algorithms that rely on sharing values or policy parameters. This shows that SS is an effective method of learning cooperative policies for collective return by sharing suggestions instead of values or policies.
In Fig. 1, SS converges faster than PS, which implies that sharing action distributions is more efficient than sharing parameters of policy networks. In Fig. 2, SS outperforms both VS and VPS in almost all the tasks. Additionally, PS shows better performance than VS and VPS, which may indicate that sharing policy information is more effective than sharing value information. Notably, SS outperforms CL in some cases. We hypothesise that in these scenarios, SS facilitates cooperation by enabling agents to encourage each other through action suggestions based on their individual interests, while CL may struggle due to exploration issues arising from a lack of successful cooperation experiences. IMR also shows competitive performance, even achieving the best results in a specific case. However, for the problem addressed in this paper, adding intrinsic rewards may not always be practical, especially in scenarios where designing appropriate intrinsic rewards is challenging.